OpenStudy (anonymous):

factor the following sum of two cubes z^3 +64

jhonyy9 (jhonyy9):

64 = what number cubed ?

jhonyy9 (jhonyy9):

so you need to know again that for sum and differents of two cubes there exist formules you dont have learned again in your schoole ?

OpenStudy (anonymous):

a^3 + b^3= (a^2 - ab + b^2)

OpenStudy (anonymous):

a^3 + b^3= (a^2 - ab + b^2)(a+b)

jhonyy9 (jhonyy9):

this right sure ?

jhonyy9 (jhonyy9):

Jenny ? what you have wrote before

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

@Jenny4218 , can you answer @jhonyy9 's question: \(\Large 64=(???)^3 \)

OpenStudy (anonymous):

can you think of a number such that when you multiply it by itself three times you'd get 64? in other words, x * x * x = 64.... what's x?

OpenStudy (anonymous):

4

jhonyy9 (jhonyy9):

yes right so than this formula will be how ?

OpenStudy (anonymous):

not sure

jhonyy9 (jhonyy9):

why not just you need replace 4 in place of b and z in place of a ok ?

OpenStudy (anonymous):

oka!

OpenStudy (anonymous):

(z+4)(z^2 -4z +4^2)

OpenStudy (anonymous):

looks good... you may want to simplify 4^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u help with this:

OpenStudy (anonymous):

10x^3 + 80

jhonyy9 (jhonyy9):

factoriz out 10 and will get 10(x^3 +8)= can you continue it ?